A stone is released from the top of the tower. It covers 24.5m distance in the last second of its journey. What is the height of the tower 

Grade:12th pass

1 Answers

Khimraj
5 years ago
Height of the tower = H
Total time taken to reach the ground = T
H = 0.5gT^2 ——(1)
H - 24.5 = 0.5g * (T - 1)^2
0.5gT^2 - 24.5 = 0.5g * (T^2 + 1 - 2T)
0.5gT^2 - 24.5 = 0.5gT^2 + 0.5g - gT
gT - 24.5 = 0.5g
10T - 24.5 = 5
10T = 29.5
T = 2.95 s
Substitute T = 2.95 s in equation (1)
H = 0.5gT^2
= 0.5 * 10 * (2.95)^2
= 43.5 m
Height of the tower is 43.5 m